# Deep Reinforcement Learning-Based Active Flow Control of an Elliptical Cylinder: Transitioning from an Elliptical Cylinder to a Circular Cylinder and a Flat Plate

### v1

@@ Abstract
This study investigates the effectiveness of Active Flow Control (AFC) technology supported by Deep Reinforcement Learning (DRL) in suppressing vortex shedding, reducing drag, and minimizing lift coefficients around elliptical cylinders. The aspect ratio ($Ar$) of the elliptical cylinder transitions from an ellipsoid ($Ar=2.0$) to a circular shape ($Ar=1.0$), and ultimately to a flat plate ($Ar=0$). We utilize the Proximal Policy Optimization (PPO) algorithm to precisely control the mass flow rates of synthetic jets located on both the upper and lower surfaces of the cylinder. The focus is on optimizing fluid dynamics and examining the scalability of these control techniques across different geometric configurations. Our research findings indicate that, for elliptical cylinders with aspect ratios between 1.75 and 0.75, the reduction in drag coefficient ranges from 0.9% to 15.7%, and the reduction in lift coefficient ranges from 95.2% to 99.7%. Notably, the DRL-based control strategy not only significantly reduces lift and drag, but also completely suppresses vortex shedding while using less than 1% of external excitation energy, demonstrating its efficiency and energy-saving capabilities. Additionally, for aspect ratios from 0.5 to 0, the reduction in drag coefficient ranges from 26.9% to 43.6%, and the reduction in lift coefficient from 50.2% to 68.0%. This reflects the control strategy’s significant reduction in both drag and lift coefficients, while also alleviating vortex shedding. This underscores the adaptability and potential of DRL-based AFC in managing complex fluid dynamics across diverse elliptical cylinder configurations.
@@ I Introduction
The development of active flow control technology has progressed from theoretical exploration to practical application, gradually becoming an important branch within the domains of fluid mechanics and engineeringCollis and Joslin (2004); Jahanmiri (2010); Awada, Younes, and Ilinca (2021). This technology aims to improve flow characteristics through active intervention in the flow field, such as reducing drag, controlling vortex shedding, enhancing mixing efficiency, or preventing flow separation, thereby enhancing system performance and efficiencyCollis and Joslin (2004); Wang et al. (2018). Initially, research in active flow control primarily focused on theoretical analysis and simple experiments, exploring how to influence fluid flow with external inputs like jets, vibrations, or magnetic fieldsJahanmiri (2010); Wang et al. (2018); Awada, Younes, and Ilinca (2021). With the advancement of Computational Fluid Dynamics (CFD) and increased computational power, researchers have been able to gain a deeper understanding of complex flow responses through numerical simulations and develop more effective control strategiesAram et al. (2018); Jain, Yeo, and Chopra (2010); Li et al. (2022). In recent years, breakthroughs in machine learning and deep learning technologies have led to revolutionary progress in the field of active flow controlBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Ren, Hu, and Tang (2020). Researchers have begun using intelligent algorithms, such as Deep Reinforcement Learning (DRL), to automatically optimize control strategies, achieving precise control over highly nonlinear and variable flow fieldsBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Garnier et al. (2021). These methods have not only improved the accuracy and real-time performance of control effects but also significantly expanded the application scope of active flow control technology, including but not limited to aerospace, automotive, energy, and environmental engineering sectorsVignon, Rabault, and Vinuesa (2023); Rabault et al. (2020); Garnier et al. (2021); Razdan and Shah (2022).
@@
DRL is an avant-garde methodology that amalgamates the principles of deep learning with reinforcement learning to optimize decision-making through interaction with an environmentArulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2022a). Deep learning excels in autonomously extracting intricate features from voluminous datasets, while reinforcement learning concentrates on how to make optimal decisions based on these features, guiding the learning process through rewardsJaniesch, Zschech, and Heinrich (2021); LeCun, Bengio, and Hinton (2015). DRL leverages deep neural networks to represent the policies or value functions within reinforcement learning, thus empowering machines to tackle high-dimensional state spaces and complex decision-making challenges that were previously intractableArulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2022a). Notable for its robust non-linear modeling capabilities and adaptable learning mechanisms, DRL is capable of directly learning complex strategies from raw sensory data without the need for manually engineered featuresDargazany (2021). Moreover, DRL is inherently self-improving, continuously refining its strategies through ongoing interactions with the environmentWang et al. (2022a). This capacity for self-learning and adaptation bestows a significant advantage upon DRL, especially in resolving decision-making problems within dynamic settingsFrançois-Lavet et al. (2018); Wang et al. (2022a). The applications of DRL are extensive, encompassing domains such as gaming, autonomous vehicles, robotic control, natural language processing, and computer vision. In the gaming industry, breakthroughs like AlphaGo’s victory over world champions underscore DRL’s potentialGranter, Beck, and Papke (2017); in autonomous driving, DRL is employed for optimizing route planning and decision-makingSallab et al. (2017); in robotics, it enables robots to learn complex physical tasksZhao, Queralta, and Westerlund (2020). Additionally, DRL is being applied in scientific research areas like fluid mechanics and industrial optimization, showcasing its vast potential in managing complex systemsBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015).
In recent years, the application of DRL within the realm of fluid dynamics has seen a marked increase, with a primary focus on active flow control and shape optimizationBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Garnier et al. (2021). Active flow control technology is a key challenge in the fields of fluid mechanics and engineering, including goals such as drag reduction, prevention of flow separation, control of vortex shedding, and enhancement of mixing efficiencyWang et al. (2018). Traditional control methodologies often rely on a priori knowledge and empirical rules, whereas DRL offers a novel pathway for autonomously learning optimal control strategies from dataCollis and Joslin (2004); Jahanmiri (2010). In the context of drag reduction applications, DRL is utilized to autonomously adjust the operation of control devices (e.g., jets, vibratory mechanisms) to minimize the fluidic drag around objects such as cylindrical bodies or car modelsGarnier et al. (2021); Razdan and Shah (2022). For vortex shedding control, DRL is capable of learning how to dynamically adjust control parameters to stabilize the flow field and reduce vibrations and noise caused by vortex sheddingVignon, Rabault, and Vinuesa (2023); Rabault et al. (2020). In the domain of shape optimization, DRL facilitates the automatic discovery and application of effective shape adjustments to achieve optimization goals by intelligently exploring the design space and learning optimal design strategiesBhola et al. (2023). With the continued expansion of computational resources and advancements in algorithms, the application of DRL in fluid dynamics is set to broaden further.
Fluid mechanics problems are characterized by a high degree of nonlinearity and multiscalar nature, involving high-dimensional state spaces and complex system dynamicsPozorski and Wacławczyk (2020); Reichstein et al. (2019). These challenges are often intractable for traditional mathematical models, demanding sophisticated control strategies, particularly when addressing phenomena such as unstable flows, turbulence, and multiphase flows, which require consideration of a wide range of spatiotemporal scalesSubramaniam (2020). The suitability of DRL methods for these kinds of fluid mechanics issues primarily rests on several key features: Firstly, DRL can handle the high-dimensional and complex state spaces intrinsic to fluid dynamics, effectively capturing the intricate dynamics presentXie et al. (2023). Through the function approximation capabilities of deep neural networks, DRL is adept at processing fluid mechanics systems with numerous variables and parameters, discerning patterns and regularities hidden within dataZhao et al. (2024). Secondly, DRL’s interactive and adaptive learning approach allows it to self-tune and optimize strategies in response to dynamic and uncertain fluid environments. Traditional control methods often struggle with such dynamicsBhansali et al. (2023); however, DRL models, through their interaction with the environment and trial-and-error processes, can independently learn and refine strategies to meet diverse fluidic conditions and task requirementsHusen, Chaudary, and Ahmad (2022). Moreover, the long-term reward mechanism intrinsic to DRL allows for the consideration of future impacts when formulating strategies. Control and optimization in fluid mechanics often necessitate foresight into long-term effectsVinuesa et al. (2022). DRL’s focus on long-term rewards enables models to balance short-term gains with future outcomes, thereby better addressing control and optimization challenges within fluid mechanicsViquerat et al. (2022); Vinuesa (2024). In conclusion, the attributes of DRL render it highly suitable for addressing the control and optimization challenges in fluid mechanics.
Rabault et al.Rabault et al. (2019) pioneered the application of advanced DRL algorithms in active flow control. The author uses the DRL agent to control the mass flow rate of the synthetic jets on the upper and lower sides of the cylinder, effectively reducing the resistance around the cylinder by 8% and mitigating the detachment of the rear vortex. Rabault et al.; Tang et al.; Heess et al.; Ren, Rabault, and Tang; Jia and XuRabault et al. (2019); Tang et al. (2020); Heess et al. (2017); Ren, Rabault, and Tang (2021); Jia and Xu (2024a) applied the Proximal Policy Optimization (PPO) algorithm to active flow control (AFC) in the flow around a cylinder. The training results consistently showed a reduction in drag of 8% around the cylinder. Li and ZhangLi and Zhang (2022) investigated the flow stability and sensitivity in confined cylinder wakes under various blockage ratios. They used the analysis of flow stability and structure sensitivity to guide the manipulation of the flow field by DRL agents. Wang et al.Wang et al. (2022b) utilized an AFC method based on the PPO algorithm to reduce the drag on a cylindrical body by approximately 8%, and applied the Soft Actor-Critic (SAC) algorithm to decrease the drag around a square cylinder by about 13.6%. Jia and XuJia and Xu (2024b) employed the SAC algorithm within a DRL framework to achieve AFC around a square cylinder, effectively modulating the flow dynamics across varying Reynolds numbers. Wang et al.Wang et al. (2022c) used DRL methods for active control of flow through elliptical cylinders. Their training results prove that DRL can successfully learn active control strategies under four angles of attack and various long and short axis ratios. Fan et al.Fan et al. (2020) was the first to apply DRL to active flow control experiments, optimizing power efficiency by adjusting the rotational speed of two small cylinders behind the main cylinder. Ren, Wang, and TangRen, Wang, and Tang (2021) proposed a novel active flow control strategy to hide the hydrodynamic characteristics of bluff bodies, including strong shear forces and periodic shedding vortices, to avoid predators.
The application of AFC strategies based on DRL has been widely validated in various bluff body flow scenarios for their effectiveness in controlling vortex shedding and reducing drag. Among these, studies on flow around constrained cylinders under different blockage ratios represent the most thoroughly investigated areaRabault et al. (2019); Tang et al. (2020); Li and Zhang (2022); Wang et al. (2022b); Jia and Xu (2024a). By contrast, research related to flow around elliptic cylinders has seen relatively less exploration within this domain. Due to their unique shape, elliptic cylinders influence the development of separation points and boundary layers differently during flow, potentially leading to Hopf bifurcation and periodic vortex shedding at lower Reynolds numbers. These flow characteristics offer opportunities in aerospace, marine engineering, energy engineering, and modern architectural design for improving lift-to-drag ratios, reducing vortex-induced vibration, enhancing energy efficiency, and increasing structural stability. Therefore, an in-depth investigation of the flow characteristics around elliptic cylinders not only enhances our understanding of fluid dynamics but also promotes performance optimization and innovation in engineering design.
In summary,despite significant advancements in controlling flow around circular cylinders through DRL algorithms, the study of flow around elliptical cylinders presents a vast research landscape due to its more complex flow characteristics and control challenges. AFC based on DRL algorithms introduces a novel methodology that leverages learning from flow field information to intelligently adjust control strategies, aiming at reducing drag and suppressing vortex shedding oscillations. This research is dedicated to examining the impact of various $Ar$ of elliptical cylinders on flow stability and sensitivity to structural parameters. Building upon this, it explores the application of DRL-based AFC in managing flow around elliptical cylinders by utilizing the derived physical insights. By establishing a robust reinforcement learning environment and defining clear reward mechanisms, this study aims to identify and implement effective flow control strategies. The aim of this study is to enhance the flow stability around an elliptical cylinder using AFC technology based on DRL algorithm, serving as a reference for flow control around similarly complex-shaped objects.
@@
The structure of this paper is organized as follows: Section II introduces the problem description and provides a comprehensive overview of Deep DRL and its framework for application in AFC. Section Section III presents the main results and analysis. Finally, Section IV concludes the paper by summarizing the key findings and discussing their implications.
@@ II.1 Numerical simulation
The present study investigates the flow around two-dimensional elliptical cylinders with their major axis oriented perpendicular to the flow direction, as depicted in Fig. 1(a). The major axis of the constrained elliptical cylinder (denoted 2a) remains fixed, while the shape of the elliptical cylinder is changed by changing the value of 2b. The Reynolds number is calculated based on the length of the major axis, formulated as $Re=U(2a)/\nu$, where $U$ represents the velocity of the incoming uniform flow and $\nu$ denotes the kinematic viscosity. The aspect ratio ($Ar$) is defined as the ratio of the semi-minor axis ($b$) to the semi-major axis ($a$), expressed as $Ar=b/a$. For ease of comparison with circular cylinders, we define $2a$ as $D$, representing the diameter of an equivalent circular cylinder. Consequently, an $Ar$ of 0 designates a flat plate normal to the flow, while an $Ar$ of 1 denotes a circular cylinder. This study investigates a series of variations ranging from an elliptical cylinder $(Ar=2.0)$ to a circular cylinder $(Ar=1.0)$, and further to a flat plate $(Ar=0)$, as shown in Fig. 1(c).
Figure 1: (a) Definition of the problem, including relevant dimensions. (b) Establishment of boundary conditions. (c) This study examines an elliptical cylinder with varying aspect ratios: a normal flat plate corresponding to a cylinder with an aspect ratio of 0, a special case of a cylinder with an aspect ratio of 1, and an elliptical cylinder with an aspect ratio of 2.
Computational domain and coordinate system
An elliptical cylinder is placed within a computational domain bounded by a rectangular perimeter, where the diameter of the elliptical cylinder is defined as D, and the major axis length (2a) of the elliptical column is equivalent to D. The computational domain extends from -2D in front of the cylinder or elliptical column to 20D behind it in the flow direction, which is the direction of fluid flow. In the cross-flow direction, perpendicular to the flow direction, the domain ranges from -2D to 2.1D. Within this computational domain, a Cartesian coordinate system is established with the center of the cylinder or elliptical column as the origin O, where the direction of flow is defined along the x-axis, and the axis perpendicular to the flow direction is the y-axis. The configuration of the synthetic jets is depicted in Fig. 1(b), specifically positioned at two directions on the cylindrical or elliptical column, $\theta_{1}=90^{\circ}$ and $\theta_{2}=270^{\circ}$, representing symmetrical locations of the jet on either side of the object’s surface. The width of the jet opening is fixed at $\omega=10^{\circ}$.
@@
Let $\Omega\subset\mathbb{R}^{nd}$ and $(0,T)$ be the spatial and temporal domains, respectively, where $nd$ is the number of space dimensions, and let $\Gamma$ denote the boundary of $\Omega$. The spatial and temporal coordinates are denoted by $x$ and $t$. This work investigates the dynamics of incompressible viscous fluid flow. The Navier–Stokes equations governing incompressible fluid flow are
@@
$\nabla\cdot\mathbf{u}=0\quad\text{on}\quad\Omega\times(0,T).$ (1b)
Here, $\bm{u}=(U,V)$ represents the velocity vector, where $U$ and $V$ are the velocity components along the $x$- and $y$-axes, respectively. $p$ is the thermodynamic pressure. The $Re$ is calculated based on the length of the major axis, formulated as $Re=U(2a)/\nu$, where $U$ represents the velocity of the incoming uniform flow and $\nu$ denotes the kinematic viscosity.
@@
The upstream boundary of the computational domain, designated as $\Gamma_{\text{in}}$, is characterized by a parabolic velocity profile oriented in the $x$-axis, where the boundary conditions are specified such that the streamwise velocity component $U$ is equal to the mean velocity $U$ with a value of 1, and the transverse velocity component $V$ is set to zero. On the downstream boundary, denoted as $\Gamma_{\text{out}}$, a Neumann-type boundary condition is prescribed for the velocity field, ensuring that the stress vector is nullified. On the upper and lower boundaries $\Gamma_{\text{w}}$ , a slip-wall boundary condition is employed, which stipulates that both the velocity component perpendicular to these boundaries and the tangential stress vector component along them are assigned a value of zero. The synthetic jets boundaries are denoted by $\Gamma_{i}(i=1,2)$. The velocity within the jets adheres to a parabolic distribution, with the sign of the jet velocity indicative of fluid suction or injection. The design of the synthetic jets ensures mass flow balance, as expressed by the condition $V_{\Gamma_{1}}=-V_{\Gamma_{2}}$, to maintain mass conservation in the system. Except for the jet, the non-slip solid wall boundary condition $\Gamma_{\text{w}}$ is adopted for the rest of the cylinder.
@@
Here, $U_{m}$ is the maximum velocity magnitude of the parabolic profile, and $H=4.1D$ represents the total height of the rectangular domain.
The functions $f_{Q_{i}}$ represent radial velocity profiles that simulate the suction or injection mechanisms of the jets. These profiles are selected to ensure a smooth transition to the no-slip boundary conditions on the surfaces $\Gamma_{W}$ of the cylinder. Specifically, $f_{Q_{i}}$ is defined as $A(\theta;Q_{i})(x,y)$, with the modulation relying on the angular coordinate $\theta$ as illustrated in Fig. 2(b). For the jet characterized by a width of $\omega$ and centered at $\theta_{0}$, which is positioned on the cylinder with radius $R$, the modulation is established as follows:
@@
Taking the computational domain of an elliptical cylinder with an $Ar$ of 1 as an exemplar, the domain undergoes discretization utilizing a synergistic approach that incorporates both structured (quadrilateral) and unstructured (triangular) meshes. As delineated in Fig. 2, this process results in the domain being segmented into 18,484 mesh elements. Herein, the vicinity of the elliptical cylinder is discretized with multiple layers of quadrilateral meshes, whereas the remaining regions are discretized using triangular meshes. Fig. 2(a) reveals the global discretization scheme of the computational domain, highlighting the comprehensive approach to mesh division. Fig. 2(b) offers insight into the mesh partitioning adjacent to the elliptical cylinder, with the multi-layered quadrilateral mesh around the cylinder being significantly finer compared to the triangular mesh. Fig. 2(c) illustrates the multi-layered quadrilateral mesh division around the elliptical cylinder and details the integration strategy between structured and unstructured meshes. Lastly, Fig. 2(d) provides a closer inspection of the mesh division surrounding the elliptical cylinder, emphasizing the precision of the discretization near the structure. The mesh division method for ellipses with $Ar$ other than 1 follows the same approach as that for $Ar=1$.
Figure 2: Numerical simulation of water flowing through an elliptical cylinder, close-up view of the mesh at $Ar=1$. The computational domain is discretized into 18,484 mesh elements, incorporating both triangular and quadrilateral meshes. (a) Discretization of the computational domain is viewed from a global perspective. (b) A comprehensive display of the mesh around the elliptical cylinder. (c) Presentation of the multi-layered quadrilateral meshes surrounding the elliptical cylinder. (d) A magnified view of the quadrilateral meshes in the vicinity of the elliptical cylinder.
@@
$St=\frac{f\cdot D}{U}.$ (6)
Where $f_{s}$ is the shedding frequency calculated based on the periodic evolution of the $C_{L}$. $D$ is the characteristic length of the elliptical cylinder, and $\overline{U}$ is the mean velocity of the upstream flow.

### v2

@@ Abstract
This study investigates the effectiveness of active flow control (AFC) technology supported by deep reinforcement learning (DRL) applied to flows around elliptical cylinders at $Re=100$. We vary the aspect ratio ($Ar$) of the elliptical cylinder from an ellipsoid ($Ar=2.0$) to a circular shape ($Ar=1.0$), and ultimately to a flat plate ($Ar=0$). We utilize the proximal policy optimization (PPO) algorithm to precisely control the mass flow rates of synthetic jets located on both the upper and lower surfaces of the cylinder. The control objective focuses on reducing drag, minimizing lift and suppressing vortex shedding. We examine the robustness and adaptability of DRL-based control techniques across different geometric configurations. Our research findings indicate that, for elliptical cylinders with $Ar$ between 1.75 and 0.75, the reduction in drag coefficient ranges from 0.9% to 15.7%, and the reduction in lift coefficient ranges from 95.2% to 99.7%. Notably, the DRL-based control strategy not only significantly reduces lift and drag, but also completely suppresses vortex shedding while using less than 1% of external excitation energy, demonstrating its efficiency and energy-saving capabilities. Additionally, for $Ar$ from 0.5 to 0, the reduction in drag coefficient ranges from 26.9% to 43.6%, and the reduction in lift coefficient from 50.2% to 68.0%. This reflects the control strategy’s significant reduction in both drag and lift coefficients, while also alleviating vortex shedding. Overall, our findings underscore the adaptability and potential of DRL-based AFC in controlling complex fluid dynamics across diverse geometric configurations.
@@ I INTRODUCTION
The development of active flow control (AFC) technology has progressed from theoretical exploration to practical application, gradually becoming an important branch within the domains of fluid mechanics and engineering.Collis and Joslin (2004); Jahanmiri (2010); Awada, Younes, and Ilinca (2021) This technology actively intervenes the flow field to improve flow characteristics, such as reducing drag, controlling vortex shedding, enhancing mixing efficiency, or preventing flow separation, thereby enhancing system performance and efficiency.Collis and Joslin (2004); Wang et al. (2018) In its early stages, AFC research primarily focused on theoretical analysis and simple experiments to explore the influence of external inputs, such as jets, vibrations, or magnetic fields, on fluid flow dynamics.Jahanmiri (2010); Wang et al. (2018); Awada, Younes, and Ilinca (2021) With the advancement of computational fluid dynamics (CFD) and increased computational power, researchers have been able to delve deeper into complex flow responses through numerical simulations and develop more effective control strategies.Aram et al. (2018); Jain, Yeo, and Chopra (2010); Li et al. (2022) However, as fluid problems are typically characterized by a high degree of nonlinearity and multiscale nature, involving high-dimensional state spaces and complex system dynamicsPozorski and Wacławczyk (2020); Reichstein et al. (2019), many engineering problems are often intractable using existing models. Sophisticated control strategies that require consideration of a wide range of spatiotemporal scales are indispensable when addressing phenomena such as unstable flows, turbulence, and multiphase flows.Subramaniam (2020)
@@
Recently, the field of AFC has witnessed revolutionary progress, thanks to breakthroughs in machine learning and deep learning technologies.Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Ren, Hu, and Tang (2020) Deep reinforcement learning (DRL) is an avant-garde technique that combines the principles of deep learning with reinforcement learning to optimize decision-making through interactions with an environment.Arulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2020) Deep learning excels at autonomously extracting intricate features from voluminous datasets, while reinforcement learning concentrates on how to make optimal decisions based on these features by guiding the learning process through rewards.Janiesch, Zschech, and Heinrich (2021); LeCun, Bengio, and Hinton (2015) By utilizing deep neural networks to represent policies or value functions within reinforcement learning, DRL enables machines to tackle complex decision-making challenges and high-dimensional state spaces that were previously considered intractable.Arulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2020) Renowned for its robust non-linear modeling capabilities and adaptive learning mechanisms, DRL has the capability to directly learn complex strategies from raw sensory data without relying on manually engineered features.Dargazany (2021) Furthermore, DRL is inherently self-improving, continuously refining its strategies through ongoing interactions with the environment.Wang et al. (2020) The self-learning and adaptive nature of DRL grants a significant advantage, particularly in resolving decision-making problems in dynamic settings.François-Lavet et al. (2018); Wang et al. (2020) The applications of DRL are vast and encompass various domains, including gaming, autonomous vehicles, robotic control, natural language processing, and computer vision. In the gaming industry, distinguished achievements like AlphaGo’s victory over world champions highlight the potential of DRL.Granter, Beck, and Papke (2017) In the realm of autonomous driving, DRL is employed to optimize route planning and decision-making.Sallab et al. (2017) In the field of robotics, it enables robots to learn complex physical tasks.Zhao, Queralta, and Westerlund (2020) Overall, DRL offers the potential to develop adaptive and intelligent control strategies that can adapt to changing environments and optimize performance in a wide range of applications.
With progressive advancement in the DRL technology, there has been a significant increase in the application of DRL in the field of fluid dynamics as well.Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Garnier et al. (2021); Razdan and Shah (2022) DRL methods are highly suited to handle the high-dimensional and complex state spaces intrinsic to fluid dynamics.Xie et al. (2023) Through the function approximation capabilities of deep neural networks, DRL is adept at processing fluid systems and discerning flow patterns and structures.Zhao et al. (2024) DRL models, through their interaction with the environment and trial-and-error processes, can independently learn and refine strategies to meet diverse fluid conditions.Husen, Chaudary, and Ahmad (2022); Collis and Joslin (2004); Jahanmiri (2010) For instance, DRL has been applied to shape optimization of fluid systems and enabled automatic discovery and application of effective shape adjustments to achieve optimization goals by intelligently exploring the design space and learning optimal strategies.Bhola et al. (2023) Moreover, the long-term reward mechanism intrinsic to DRL allows for the consideration of future impacts when formulating strategies. This unique feature has allowed DRL to shine in many AFC problems where DRL’s focus on long-term rewards enables control models to balance short-term gains with future outcomes, thereby better addressing control and optimization challenges within fluid mechanics.Vinuesa et al. (2022); Vignon, Rabault, and Vinuesa (2023); Viquerat et al. (2022); Vinuesa (2024)
In the academic literature, flow control around blunt bodies using DRL has attracted considerable interest due to its pivotal role in reducing drag, a key aspect of fluid dynamics research. This pursuit carries several important implications, including enhanced energy efficiency, improved performance, noise and vibration reduction, and industrial process optimization. As flow around blunt body tends to involve strong non-linearities such as flow separation and recirculation, traditional control methods often struggle in achieving control objectives whereas DRL models have shown promising results. Pioneered by Rabault et al. Rabault et al. (2019), various researchers such as Tang et al. Tang et al. (2020), Heess et al. Heess et al. (2017), Ren, Rabault, and Tang Ren, Rabault, and Tang (2021), and Wang and Xu Wang and Xu (2024) applied the Proximal Policy Optimization (PPO) algorithm to AFC in the flow around a circular cylinder. Consistently, they achieved an approximate 8% reduction in drag in their validation experiments. Furthermore, Li and ZhangLi and Zhang (2022) ,Wang et al.Wang et al. (2022a), Jia and XuJia and Xu (2024), and Wang et al.Wang et al. (2022b) respectively conducted flow control studies on confined circular cylinders and square cylinders using DRL. In addition, Fan et al.Fan et al. (2020) initiated the application of DRL to AFC experiments, optimizing power efficiency by adjusting the rotational speed of auxiliary circular cylinders positioned behind another circular cylinder. Ren, Wang, and TangRen, Wang, and Tang (2021) proposed an AFC strategy to conceal the hydrodynamic characteristics of circular cylinders, such as strong shear forces and periodic shedding vortices. Studies on flow around constrained circular cylinders under different blockage ratios have also received extensive attention in the field Rabault et al. (2019); Tang et al. (2020); Li and Zhang (2022); Wang et al. (2022a); Wang and Xu (2024). The existing studies have successfully demonstrated the effectiveness of using DRL for flow control, particularly in actively manipulating the flow around circular cylinders and square cylinders.
Although active flow control strategies for both circular and square cylinders are pivotal in various engineering applications for managing aerodynamic and hydrodynamic forcesVignon, Rabault, and Vinuesa (2023); Rabault et al. (2020); Garnier et al. (2021); Razdan and Shah (2022), their extension to more intricate, non-ideal geometries poses significant challenges. The effectiveness of DRL strategies in such contexts remains largely unexplored, marked by uncertainties regarding their capacity to generate physically reasonable and adaptable control measures. Current literature indicates a distinct lack of comprehensive studies on the variability of DRL-based flow control strategies across different geometries, such as from streamlined bodies to various bluff bodies. This research proposes to bridge these gaps by focusing on the flow dynamics around parameterized ellipsoids, examining how variations in the aspect ratio influence flow fields and their corresponding DRL-driven control strategies. By extending the investigation from elongated ellipsoids to circular cylinders and subsequently normal flat plates, this study aims to develop a deeper understanding of the physical validity of DRL approaches and their adaptability across varying body geometries.
@@
This work extends the studies conducted by Wang and XuWang and Xu (2024); Jia and Xu (2024) by applying the DRL-based AFC approach to elliptical cylinders, which present more complex flow characteristics and additional control challenges. The objective is to develop effective control strategies to reduce drag, improve flow stability, and enhance overall performance. The structure of this paper is organized as follows: Sec. II introduces the problem description and provides a comprehensive overview of DRL and its framework for application in AFC. Sec. III presents the main results and analysis. Finally, Sec. IV concludes the paper by summarizing the key findings and discussing their implications.
@@ II.1 Numerical simulation
[…]
Figure 1: (a) Definition of the problem, including the dimensions of the computational domain and the establishment of a Cartesian coordinate system. (b) Establishment of boundary conditions; note that the dimensions of the elliptical cylinder and the computational domain in the figure are not to scale, focusing solely on the delineation of boundary conditions. (c) This study investigates elliptical cylinders with different aspect ratio: a normal flat plate corresponding to $Ar=0$, a circular cylinder with $Ar=1$, and an elliptical cylinder with $Ar=2$. The dimension $2a$ is fixed as $D$, and the $Ar$ is varied by changing the length of 2b.
@@
Let $\Omega\subset\mathbb{R}^{nd}$ and $(0,T)$ be the spatial and temporal domains, respectively, where $nd$ is the number of space dimensions, and let $\Gamma$ denote the boundary of $\Omega$. The Navier-Stokes equations that describe the dynamics of this incompressible viscous fluid flow are given by:
@@
$\nabla\cdot\mathbf{u}=0\quad\text{on}\quad\Omega\times(0,T),$ (1b)
where $\mathbf{u}$, $t$ and $p$ are the velocity vector, time, and pressure, respectively. In this study, the $Re$ is set to 100.
@@
The upstream boundary of the computational domain, designated as $\Gamma_{\text{in}}$, is characterized by a parabolic velocity profile oriented in the $x$-axis, where the boundary conditions are specified such that the transverse velocity component is set to zero, and the velocity along the stream direction shows a parabolic profile distribution. On the downstream boundary, denoted as $\Gamma_{\text{out}}$, a Neumann-type boundary condition is prescribed for the velocity field, ensuring that the stress vector is nullified. On the upper and lower boundaries $\Gamma_{\text{w}}$ , a slip-wall boundary condition is employed, which stipulates that both the velocity component perpendicular to these boundaries and the tangential stress vector component along them are assigned a value of zero. The synthetic jets boundaries are denoted by $\Gamma_{i}(i=1,2)$. The velocity within the jets adheres to a parabolic distribution, with the sign of the jet velocity indicative of fluid suction or injection. The design of the synthetic jets ensures mass flow balance, as expressed by the condition $V_{\Gamma_{1}}=-V_{\Gamma_{2}}$, to maintain mass conservation in the system. Except for the synthetic jets, the non-slip solid wall boundary condition $\Gamma_{\text{w}}$ is adopted for the rest of the cylinder.
@@
Here, $U_{m}$ is the maximum velocity magnitude of the parabolic profile, and $H=4.1D$ represents the total height of the rectangular domain. The mean velocity magnitude is $\overline{U}=2U_{m}/3=1$ and is used to define the Reynolds number $Re=\overline{U}D/\nu$.
The functions $f_{Q_{i}}$ represent radial velocity profiles that simulate the suction or injection mechanisms of the jets. These profiles are selected to ensure a smooth transition to the no-slip boundary conditions on the surfaces $\Gamma_{W}$ of the cylinder. Specifically, $f_{Q_{i}}$ is defined as $A(\theta;Q_{i})(x,y)$, with the modulation relying on the angular coordinate $\theta$ as illustrated in Fig. 2. For the jet characterized by a width of $\omega$ and centered at $\theta_{0}$, which is positioned on the cylinder with radius $R$, the modulation is established as follows:
@@
Considering an elliptical cylinder computational domain with $Ar=1$, the domain is discretized employing a hybrid meshing approach that integrates both structured (quadrilateral) and unstructured (triangular) meshes. As delineated in Fig. 2, the computational domain is divided into 18,484 mesh cells. Specifically, the area surrounding the elliptical cylinder is discretized using a multi-layer quadrilateral mesh, while the remaining regions are discretized with triangular meshes. Fig.2(a) displays the global discretization scheme for the computational domain, highlighting the comprehensive approach to mesh division. Fig.2(b) provides a detailed view of the mesh partitioning near the elliptical cylinder, where the multi-layered quadrilateral meshes surrounding the cylinder are notably finer compared to the triangular meshes. Fig.2(c) illustrates the division of multi-layered quadrilateral meshes around the elliptical cylinder, emphasizing the precision of discretization near the structure. Finally, fig.2(d) showcases the mesh division around the elliptical cylinder and the integration strategy between quadrilateral and triangular meshes. The method for mesh division for ellipses with $Ar$ values other than 1 follows the same approach as that for $Ar=1$, and is not separately demonstrated.
Figure 2: For $Ar=1$, the computational domain is discretized into 18,484 mesh cells, including both triangular and quadrilateral meshes. (a) Discretization of the computational domain is viewed from a global perspective. (b) A comprehensive display of the mesh around the elliptical cylinder. (c) Presentation of the multi-layered quadrilateral meshes surrounding the elliptical cylinder. (d) A magnified view of the quadrilateral meshes in the vicinity of the elliptical cylinder.
@@
$St=\frac{f_{s}\cdot D}{\overline{U}}$ (6)
Where $f_{s}$ is the shedding frequency calculated based on the periodic evolution of the $C_{L}$. $D$ is the characteristic length of the elliptical cylinder, and $\overline{U}=1$.